### v1

@@ 3.2 Confounding Variables and Simpson’s Paradox
However, this correlation may be confounded by operational modes $Z_{\text{mode}}(t)$ (normal flow, backwash, cleaning), leading to Simpson’s Paradox [36] (Eq. 5):
@@
(a) Observational analysis
(b) Observational analysis
(c) Observational analysis
Figure 2: Three Levels of Causal Analysis: (a) Observational analysis results, (b) Intermediate analysis, and (c) Final analysis.
3.3 Proposed Causal Digital Twin Framework
To solve these limitations, we propose a new Causal Digital Twin (CDT) framework that models causal relations explicitly using Structural Causal Models (SCMs). We represent the SWaT system as a causal graph (Eq. 6):
@@ 3.3 Proposed Causal Digital Twin Framework
The causal digital twin function is defined as (Eq. 8):
@@ 3.4 Research Objectives
This work aims to develop a novel causal reasoning framework for digital twins in cyber-physical security, specifically: (1) construct causal graphs from SWaT operational and attack data using causal discovery algorithms, (2) develop identification strategies for causal effects under partial observability, (3) implement do-calculus for interventional queries in real-time ICS monitoring, (4) validate counterfactual reasoning for attack attribution and defense planning, and (5) demonstrate superiority over correlation-based approaches in 41 documented SWaT attacks.
@@ 4 Methodology
This section shows our new Causal Digital Twin (CDT) framework for cyber-physical security in the SWaT system. The method combines causal inference theory with digital twin modeling using five main components.
@@ 4.1 Causal Graph Discovery
We employ the PC algorithm enhanced with time-series considerations [32, 31]. The algorithm proceeds through skeleton discovery using conditional independence tests (Eq. 11):
@@
$\rho_{ij|\mathcal{S}}^{(\ell)}=\frac{\text{Cov}(X_{i}(t),X_{j}(t-\ell)\mid\mathcal{S})}{\sqrt{\text{Var}(X_{i}(t)\mid\mathcal{S})\,\text{Var}(X_{j}(t-\ell)\mid\mathcal{S})}}$ (12)
Orientation rules apply causal ordering constraints: temporal precedence ($X_{i}(t-\ell)\rightarrow X_{j}(t)$ for $\ell>0$), physical constraints (flow sensors $\rightarrow$ level sensors), and control logic (controller outputs $\rightarrow$ actuator states).
@@ 4.2 Structural Causal Model Estimation
4.3 Interventional Digital Twin Construction
To compute interventional distributions $P(Y|\text{do}(X=x))$, we implement Pearl’s do-calculus rules. For identifiable causal effects, we apply adjustment formulas  [34]:
@@ 4.4 Causal Anomaly Detection and Root Cause Analysis
$\text{MCAI}(t)=\sum_{i=1}^{n}\alpha_{i}\cdot\text{Score}_{\text{causal}}(V_{i},t)$ (19)
@@ 4.8 Algorithm Implementation
Algorithm 1 Causal Digital Twin Framework for Cyber-Physical Security
@@
3:// Phase 1: Causal Structure Discovery
@@
7:// Phase 2: Structural Causal Model Estimation
8:for $i=1$ to $n$ do
9:  $\text{PA}(V_{i})\leftarrow$ GetParents($V_{i}$, $\mathcal{G}$)
10:  $\hat{\boldsymbol{\theta}}_{i}\leftarrow$ EstimateParameters($V_{i}$, $\text{PA}(V_{i})$, MLE)
11:end for
@@
13:// Phase 3: Model Validation
14:ValidateModel($\mathcal{M}$, CrossValidation, InterventionalTests)
15:// Phase 4: Real-time Causal Monitoring
16:for each new time step $t>T$ do
17:  for $i=1$ to $n$ do
@@
19:   $\text{Score}_{\text{causal}}(V_{i},t)\leftarrow\frac{|V_{i}(t)-\hat{V}_{i}(t)|}{\sigma_{V_{i}|\text{do}(\text{PA})}}$
20:  end for
21:  $\text{MCAI}(t)\leftarrow\sum_{i=1}^{n}\alpha_{i}\cdot\text{Score}_{\text{causal}}(V_{i},t)$
22:  if $\text{MCAI}(t)>\theta_{\text{anomaly}}$ then
23:   for $i=1$ to $n$ do
24:    $\text{CE}_{i}\leftarrow$ ComputeCausalEffect($V_{i}$, MCAI, $\mathcal{M}$)
25:    $\phi_{i}\leftarrow$ ComputeShapleyValue($V_{i}$, $\mathcal{M}$)
26:   end for
27:   $\text{RootCauses}\leftarrow$ RankByImportance($\{\text{CE}_{i},\phi_{i}\}$)
28:   for each defense mechanism $d\in\mathcal{D}$ do
29:    $P_{\text{counterfactual}}\leftarrow$ ThreeStepCounterfactual($d$, Evidence(t)(t), $\mathcal{M}$)
30:   end for
31:   return $\langle\text{RootCauses},\text{DefenseRecommendations}\rangle$
32:  end if
33:end for
@@ 5 Results
This section presents a detailed experimental evaluation of our Causal Digital Twin (CDT) framework, showing its effectiveness across multiple industrial control system datasets. We validate the framework’s main contributions through experiments on causal structure discovery, anomaly detection, root cause analysis, and deployment considerations.
@@ 5.1 Experimental Framework and Validation Methodology
We evaluate the CDT framework using three industrial control system datasets to demonstrate generalizability across different environments. The primary dataset is the Secure Water Treatment (SWaT) testbed, with 7 days of normal operation (496,800 samples at 1Hz) and 4 days containing 41 documented cyber-physical attacks across 51 variables (25 sensors, 26 actuators) in 6 process stages. Additionally, we use the Water Distribution (WADI) dataset with 14 days of normal operation and 15 attack scenarios across 123 variables, and the Hardware-in-the-loop Augmented ICS (HAI) dataset with 4 days of normal and 4 days of attack data across 78 variables.
Experiments run on Google Colab Pro with Tesla V100 GPU (16GB VRAM) using Python, causal-learn v0.1.3.1, networkx v3.1, and related libraries. All experiments use fixed random seeds and 10-fold repetition for statistical validation. Full code is available to ensure reproducibility. We compare our CDT framework against seven state-of-the-art baseline methods including CNN-1D, LSTM-ED, InterFusion, RANSynCoder, PbNN (Physics-based Neural Network), Causality-Inspired approaches, and ICS-CAD. Evaluation metrics include detection performance (Precision, Recall, F1-Score, AUC-ROC), root cause analysis accuracy (Top-k accuracy for $k=1,3,5$ and Mean Reciprocal Rank), counterfactual analysis (Intervention Prediction Error and Counterfactual Accuracy), and computational efficiency.
@@ 5.2 Causal Discovery and Anomaly Detection Performance
The causal structure discovery successfully produced interpretable causal graphs for all datasets. We check structural accuracy using three complementary methods: (1) Physical constraint compliance, which is $90.8\%\pm 3.2\%$ for SWaT, $87.4\%\pm 4.1\%$ for WADI, and $89.3\%\pm 2.9\%$ for HAI, measuring how well graphs follow known physical laws (for example, flow $\to$ pressure, level $\to$ pump activation); (2) Temporal consistency validation, showing $87.3\%\pm 2.8\%$ compliance with causal precedence constraints; and (3) Synthetic data validation, achieving structural Hamming distance $0.12\pm 0.03$ on generated SWaT-like systems with known causal structure.
@@
Table 7: Summary of F1-Score Performance Across All Datasets
Method SWaT WADI HAI
CNN-1D $0.841\pm 0.025$ $0.776\pm 0.035$ $0.801\pm 0.031$
LSTM-ED $0.862\pm 0.023$ $0.799\pm 0.033$ $0.825\pm 0.029$
InterFusion $0.907\pm 0.018$ $0.850\pm 0.028$ $0.872\pm 0.025$
PbNN $0.884\pm 0.021$ $0.825\pm 0.030$ $0.847\pm 0.027$
Causality-Inspired $0.926\pm 0.016$ $0.878\pm 0.025$ $0.900\pm 0.022$
ICS-CAD $0.928\pm 0.017$ $0.862\pm 0.027$ $0.888\pm 0.023$
@@ 5.3 Root Cause Analysis and Counterfactual Reasoning
Note: Top-1, Top-3, Top-5 = Top-k Accuracy, MRR = Mean Reciprocal Rank, AT = Attribution Time
@@ 5.4 Computational Performance and Deployment Scalability
Computational performance analysis shows the CDT framework is feasible for industrial deployment at different system scales. Inference latency scales linearly $\mathcal{O}(n)$, discovery time cubically $\mathcal{O}(n^{3})$, and memory usage quadratically $\mathcal{O}(n^{2})$, remaining real-time feasible for systems up to about 150 variables. Table 12 presents the computational breakdown on Google Colab Pro: causal discovery requires 23.4 minutes offline with 2.8 GB memory usage for the 51-variable SWaT system, while real-time components achieve 3.2 ms inference for anomaly detection and 12.7 ms for root cause analysis.
@@ 5.5 Statistical Validation and Ablation Analysis
The experimental validation shows that our Causal Digital Twin framework is an important improvement in cyber-physical security. It achieves statistically significant improvements in detection performance, attribution accuracy, and counterfactual reasoning, while keeping computational feasibility for industrial deployment. The comprehensive tests across different datasets, attack scenarios, and deployment conditions show that the framework is ready for real-world use in critical infrastructure protection.
@@ 6 Conclusion
Future work should focus on developing scalable causal discovery algorithms, improving stealthy attack detection through ensemble techniques, handling time-varying causal relationships, and creating standardized interfaces for SCADA system integration. Despite these limitations, the validated performance improvements, deployment feasibility, and interpretable results demonstrate that causal inference represents an essential advancement for protecting critical infrastructure against sophisticated cyber-physical threats, marking a paradigm shift from correlation-based to causation-aware security frameworks.
@@ Acknowledgements
Support from the University of Extremadura and the collaborating institutions is acknowledged. Thanks are also given to colleagues and project partners for useful comments during this research. Special thanks are expressed to Ambling Ingeniería y Servicios, S.L. in Extremadura for the collaboration and technical support.

### v2

@@ 3.2 Confounding Variables and Simpson’s Paradox
However, this correlation may be confounded by operational modes $Z_{\text{mode}}(t)$ (normal flow, backwash, cleaning), leading to Simpson’s Paradox Simpson (1951) (Eq. 5):
@@
(a) Level 1: Association (Observational Queries)
(b) Level 2: Intervention (Prediction of Action Effects)
(c) Level 3: Counterfactuals (What-if analysis on past events)
Figure 2: Three Levels of Causal Analysis (Pearl’s Hierarchy): (a) Association, (b) Intervention, and (c) Counterfactuals, illustrating the necessary complexity for robust cyber-physical security.
3.3 ProposedCDT Framework
To solve these limitations, we propose a new CDT framework that models causal relations explicitly using Structural Causal Models (SCMs). We represent the SWaT system as a causal graph (Eq. 6):
@@ 3.3 ProposedCDT Framework
The CDT function is defined as (Eq. 8):
@@ 3.4 Research Objectives
This work aims to develop a novel causal reasoning framework for DTs in cyber-physical security, specifically: (1) construct causal graphs from SWaT operational and attack data using causal discovery algorithms, (2) develop identification strategies for causal effects under partial observability, (3) implement do-calculus for interventional queries in real-time ICS monitoring, (4) validate counterfactual reasoning for attack attribution and defense planning, and (5) demonstrate superiority over correlation-based approaches in 41 documented SWaT attacks.
@@ 4 Methodology
This section shows our new CDT framework for cyber-physical security in the SWaT system. The method combines causal inference theory with DT modeling using five main components.
@@ 4.1 Causal Graph Discovery
To overcome the $\mathcal{O}(n^{3})$ computational complexity of centralized causal discovery algorithms, which limits scalability in large ICS, we employ a Decoupled Causal Discovery (DCD) strategy. This approach involves two phases: Candidate Parent Selection and Local Structure Discovery, transforming the complexity to $\mathcal{O}(n\cdot k^{3})$ where $k\ll n$.
Phase I: Candidate Parent Selection (Filtering). For each variable $V_{i}$, we use Time-Delay Mutual Information (TDMI) to identify a highly restricted set of potential parents $\mathcal{P}_{i}\subset\mathcal{V}_{\text{aug}}$ by maximizing the information transfer across different lags $\ell\in[0,\tau]$ (Eq. 11). This phase reduces the set size from $n$ to $k\leq 10$.
$\mathcal{P}_{i}=\{V_{j}(t-\ell):\text{TDMI}(V_{j}(t-\ell)\rightarrow V_{i}(t))>\delta\}$ (11)
Phase II: Local Structure Discovery. We then apply the PC algorithm only on the restricted local set $\mathcal{V}_{i}=\{V_{i}\}\cup\mathcal{P}_{i}$, as an enhanced algorithm for time-series considerations Spirtes et al. (1993); Peters et al. (2017). The algorithm proceeds through skeleton discovery using conditional independence tests (Eq. 12) localized on $\mathcal{V}_{i}$:
@@
$\rho_{ij|\mathcal{S}}^{(\ell)}=\frac{\text{Cov}(V_{i}(t),V_{j}(t-\ell)\mid\mathcal{S})}{\sqrt{\text{Var}(V_{i}(t)\mid\mathcal{S})\,\text{Var}(V_{j}(t-\ell)\mid\mathcal{S})}}$ (13)
Orientation rules apply causal ordering constraints: temporal precedence ($V_{i}(t-\ell)\rightarrow V_{j}(t)$ for $\ell>0$), physical constraints (flow sensors $\rightarrow$ level sensors), and control logic (controller outputs $\rightarrow$ actuator states). Causal discovery accuracy depends on data quality and may produce spurious relationships in presence of unobserved confounders, requiring domain expert validation of discovered structures.
@@ 4.2 Structural Causal Model Estimation
4.3 Interventional DT Construction
To compute interventional distributions $P(Y|\text{do}(X=x))$, we implement Pearl’s do-calculus rules. For identifiable causal effects, we apply adjustment formulas  Hernán & Robins (2024):
@@ 4.4 Causal Anomaly Detection and Root Cause Analysis
$\text{Multi-Component Anomaly Index}(t)=\sum_{i=1}^{n}\alpha_{i}\cdot\text{Score}_{\text{causal}}(V_{i},t)$ (20)
@@ 4.8 Algorithm Implementation
Algorithm 1 Causal DT Framework for Cyber-Physical Security
@@
3:// Phase 1: Causal Structure Discovery (Decoupled Approach)
@@
8:// Phase 2: Structural Causal Model Estimation
9:for $i=1$ to $n$ do
10:  $\text{PA}(V_{i})\leftarrow$ GetParents($V_{i}$, $\mathcal{G}$)
11:  $\hat{\boldsymbol{\theta}}_{i}\leftarrow$ EstimateParameters($V_{i}$, $\text{PA}(V_{i})$, MLE)
12:end for
@@
14:// Phase 3: Model Validation
15:ValidateModel($\mathcal{M}$, CrossValidation, InterventionalTests)
16:// Phase 4: Real-time Causal Monitoring
17:for each new time step $t>T$ do
18:  for $i=1$ to $n$ do
[…]
20:   $\text{Score}_{\text{causal}}(V_{i},t)\leftarrow\frac{|V_{i}(t)-\hat{V}_{i}(t)|}{\sigma_{V_{i}|\text{do}(\text{PA})}}$
21:  end for
22:  $\text{MCAI}(t)\leftarrow\sum_{i=1}^{n}\alpha_{i}\cdot\text{Score}_{\text{causal}}(V_{i},t)$
23:  if $\text{MCAI}(t)>\theta_{\text{anomaly}}$ then
24:   for $i=1$ to $n$ do
25:    $\text{CE}_{i}\leftarrow$ ComputeCausalEffect($V_{i}$, MCAI, $\mathcal{M}$)
26:    $\phi_{i}\leftarrow$ ComputeShapleyValue($V_{i}$, $\mathcal{M}$)
27:   end for
28:   $\text{RootCauses}\leftarrow$ RankByImportance($\{\text{CE}_{i},\phi_{i}\}$)
29:   for each defense mechanism $d\in\mathcal{D}$ do
30:    $P_{\text{counterfactual}}\leftarrow$ ThreeStepCounterfactual($d$, Evidence(t)(t), $\mathcal{M}$)
31:   end for
32:   return $\langle\text{RootCauses},\text{DefenseRecommendations}\rangle$
33:  end if
34:end for
@@ 5 Results
This section presents a detailed experimental evaluation of our CDT framework, showing its effectiveness across multiple industrial control system datasets. We validate the framework’s main contributions through experiments on causal structure discovery, anomaly detection, root cause analysis, and deployment considerations.
@@ 5.1 Experimental Framework and Validation Methodology
We evaluate the CDT framework using three industrial control system datasets to demonstrate generalizability across different environments. The primary dataset is the SWaT testbed, with 7 days of normal operation (496,800 samples at 1Hz) and 4 days containing 41 documented cyber-physical attacks across 51 variables (25 sensors, 26 actuators) in 6 process stages. Additionally, we use the Water Distribution (WADI) dataset with 14 days of normal operation and 15 attack scenarios across 123 variables, and the Hardware-in-the-loop Augmented ICS (HAI) dataset with 4 days of normal and 4 days of attack data across 78 variables.
Experiments run on Google Colab Pro with Tesla V100 GPU (16GB VRAM) using Python, causal-learn v0.1.3.1, networkx v3.1, and related libraries. All experiments use fixed random seeds and 10-fold repetition for statistical validation. Full code is available to ensure reproducibility. We compare our CDT framework against seven state-of-the-art baseline methods including CNN-1D, LSTM-ED, InterFusion, RANSynCoder, PbNN (Physics-based Neural Network), Causality-Inspired approaches, and ICS-CAD. Evaluation metrics include detection performance (Precision, Recall, F1-Score, AUC-ROC), root cause analysis accuracy (Top-k accuracy for $k=1,3,5$ and MRR), counterfactual analysis (Intervention Prediction Error and Counterfactual Accuracy), and computational efficiency.
@@ 5.2 Causal Discovery and Anomaly Detection Performance
The causal structure discovery successfully produced interpretable causal graphs for all datasets. We checked structural accuracy using three complementary methods: (1) Physical constraint compliance, which is $90.8\%\pm 3.2\%$ for SWaT, $87.4\%\pm 4.1\%$ for WADI, and $89.3\%\pm 2.9\%$ for HAI, measuring how well discovered edges align with validated physical relationships (for example, flow $\to$ pressure, level $\to$ pump activation) among relationships that could be verified; (2) Temporal consistency validation, showing $87.3\%\pm 2.8\%$ compliance with causal precedence constraints; and (3) Synthetic data validation, achieving structural Hamming distance $0.133\pm 0.02$ on generated SWaT-like systems with known causal structure.
@@
Table 4: Summary of F1-Score Performance Across All Datasets
Method SWaT WADI HAI
CNN-1D $0.841\pm 0.025$ $0.776\pm 0.035$ $0.801\pm 0.031$
LSTM-ED $0.862\pm 0.023$ $0.799\pm 0.033$ $0.825\pm 0.029$
InterFusion $0.907\pm 0.018$ $0.850\pm 0.028$ $0.872\pm 0.025$
PbNN $0.884\pm 0.021$ $0.825\pm 0.030$ $0.847\pm 0.027$
Causality-Inspired $0.926\pm 0.016$ $0.878\pm 0.025$ $0.900\pm 0.022$
ICS-CAD $0.928\pm 0.017$ $0.862\pm 0.027$ $0.888\pm 0.023$
@@ 5.3 Root Cause Analysis and Counterfactual Reasoning
Note: Top-1, Top-3, Top-5 = Top-k Accuracy, MRR, AT = Attribution Time
@@ 5.4 Computational Performance and Deployment Scalability
Computational performance analysis shows the CDT framework is feasible for industrial deployment at different system scales. By employing the Decoupled Causal Discovery (DCD) strategy, the initial cubic complexity $\mathcal{O}(n^{3})$ is effectively reduced to $\mathcal{O}(n\cdot k^{3})$ (where $k\leq 10$), significantly enhancing scalability. Inference latency scales linearly $\mathcal{O}(n)$, and memory usage quadratically $\mathcal{O}(n^{2})$, allowing the framework to remain feasible for systems up to 500 variables. Table 12 presents the computational breakdown on Google Colab Pro: the DCD-based causal discovery requires 23.4 minutes offline with 2.8 GB memory usage for the 51-variable SWaT system, while real-time components achieve 3.2 ms inference for anomaly detection and 12.7 ms for root cause analysis.
@@ 5.5 Statistical Validation and Ablation Analysis
The experimental validation shows that our CDT framework is an important improvement in cyber-physical security. It achieves statistically significant improvements in detection performance, attribution accuracy, and counterfactual reasoning, while keeping computational feasibility for industrial deployment. The comprehensive tests across different datasets, attack scenarios, and deployment conditions show that the framework is ready for real-world use in critical infrastructure protection.
@@ 6 Conclusion
Future work should focus on scalable causal discovery algorithms, better detection of stealthy attacks using ensembles, handling time-varying causal relationships, and creating standardized SCADA interfaces. Despite these limits, the validated performance improvements, deployment feasibility, and interpretable results show that causal inference is a key advancement for protecting critical infrastructure, marking a shift from correlation-based to causation-aware security frameworks.
@@ Acknowledgements
This initiative is carried out within the framework of the funds of the Recovery, Transformation and Resilience Plan, financed by the European Union (Next Generation) - National Cybersecurity Institute (INCIBE) in the project C107/23 “Artificial Intelligence Applied to Cybersecurity in Critical Water and Sanitation Infrastructures”. Support from the University of Extremadura, collaborating institutions, and project partners is acknowledged, with special thanks to Ambling Ingeniería y Servicios, S.L. for technical support.